The probability that ea new battery does not work is 0.04.
a) what is the probability that e new battery does work?
b)A firm prodused 5000 batteries. How many would you expect to be faulty?
1
Expert's answer
2011-06-06T07:16:24-0400
The probability that a new battery does not work is 0.04.
a)What is the probability that a new battery does work?
Solution:
The probability that a new battery works is: (1 -& the probability that it doesn’t work) = 1 - 0.04 = 0.06
Answer: 0.06
& b)A firm prodused 5000 batteries. How many would you expect to be faulty?
Solution:
Expected value = x*p(x) = 5000*0.04 = 200
Answer: 200